# Distributed Coordination for Heterogeneous Non-Terrestrial Networks

### v1

@@ IV-A Delay-tolerant Distributed Coordination
  • Cell Shaping: To cooperatively provide optimal coverage in initial access, the MADRL algorithm should be developed to select the optimal SSB codebook for each cell considering the UEs distribution and the capabilities of neighboring cells. One potential solution is to integrate the Graph Neural Network (GNN), which models heterogeneous NTNs as nodes with various features. Apart from that, the UEs associated with heterogeneous NTNs have distant transmission power capabilities, which limits their connections to upper-tier NTNs. Therefore, a lightweight model should also be deployed at the UE to make the association decision based on local characteristics including transmission power and battery.
  • Pilot Resource Allocation: To mitigate the uplink pilot interference in the heterogenous NTNs, the UE-level coordinated pilot resource allocation problem should be solved by the MADRL, where multiple cells learn to cooperatively minimize the channel correlation of UEs sharing the same pilot resource. However, UE-level allocation may increase the observation and action space significantly in MADRL, especially for higher-tier NTNs. Therefore, hierarchical MADRL can be utilized to divide the coverage into several low-correlated smaller areas first, and then focus on the pilot resource allocation within each smaller area. To reduce the communication overhead during centralized training, channel charting can be a potential solution by extracting the low-dimensional features from the statistical channels [13].
  • Multi-hop IAB: Although the training of the routing optimization algorithm is delay-tolerant, the user data transmission may have stringent latency requirements. Therefore, multi-hop transmission of duplicate user data leads to increased transmission latency and channel capacity demands. To solve the problem, the hierarchical goal-oriented semantic communication can be integrated with the MADRL to exploit the context of data and its importance to the task along with the latest latency, which extracts the most important data for transmission in each hop [14].
  • Multi-timescale Control: In the multi-timescale trajectory optimization problem, the NTNs in different tiers experience different timescales of observation, action, and reward in MADRL. Hence, a potential solution is to decouple the formulated MADRL problem based on the timescales, and then optimize the subproblems iteratively [15]. It is noted that goal-oriented semantic communication can also be utilized to quantify the importance of control signal transmission, which is closely coupled with the performance of the access links.
@@ IV-B Delay-sensitive Distributed Coordination
  • UL/DL Power Allocation: In this case, each NTN is required to deploy two heterogeneous agents to optimize the UL and DL power allocation, respectively. Specifically, the first agent detects the occurrence of CLI or RI based on the received signal and increases the uplink transmission power correspondingly. Then, the second agent optimizes the downlink transmission power based on the positions of neighboring cells to mitigate causing CLI or RI. It is noted that two agents of a single NTN work sequentially in the TDD frame structure, and same type of agents in different heterogeneous NTNs may work in an asynchronous way due to dynamic TDD.
@@ V Case Study
In this section, to validate the effectiveness of our proposed MADRL solutions in dealing with heterogeneous NTN challenges, we present the MADDPG method tailored for the user scheduling task in heterogeneous UAV networks in an emergency communication scenario. Specifically, we consider one tethered UAV (i.e., IAB donor) and four untethered UAVs (i.e., IAB node) are deployed in the disaster areas with a radius of 500m for downlink transmission, where the untethered UAVs achieve the backhaul transmission based on IAB via tethered UAV. Without loss of generality, we assume the traffic of each terrestrial user follows a Poisson process with parameter $\lambda=4$.
To maximize the downlink transmission throughput via distributed scheduling coordination, we propose a MADDPG algorithm for user scheduling based on the CTDE framework, where each UAV agent learns the optimal scheduling policy with local observations. It is noted that the observation of each agent consists of buffer status, historical SINR status towards its associated users, and the action space is the index of the associated users.
(a) Average downlink throughput during training.
(b) Downlink throughput of each UAV.
Figure 3: Average downlink transmission throughput during training and throughput of each UAV.
Fig. 3 presents the average downlink transmission throughput during training and throughput UAV. In Fig. 3 (a), we can observe that the deployment of untethered UAVs significantly improves the performance due to the enhancements to the cell-edge users. We can also see that the proposed MADDPG algorithm further achieves 37% performance gain compared with the round-robin scheduling policy. This is because the IAB links will get higher priority to be scheduled by the tethered UAV, which serves as the backhaul transmission of untethered UAVs. In Fig. 3 (b), we can observe that the downlink throughput of tethered UAV degrades around 16%, and the downlink throughput of each untethered UAV increases significantly, which validates the feasibility of MADRL for solving the challenges in heterogeneous NTNs.
@@ VI Conclusion
In this article, we first summarize the unique characteristics of different NTN platforms, and analyze their impact on the implementation of NTNs. We then focus on investigating the communication challenges for heterogeneous NTNs in individual layers, along with their potential coordinated solutions. Based on the emerging distributed algorithm, we further investigate the distributed MADRL algorithms tailed for each potential coordinated solution in heterogeneous NTNs, and analyze the CTDE deployment options according to their latency requirement. Importantly, we utilize MADDPG to solve the user scheduling optimization problem in heterogeneous UAVs-based wireless networks, which validates the effectiveness of distributed learning for heterogeneous NTNs.

### v2

@@ IV-A Delay-tolerant Distributed Coordination
  • Cell Shaping: To cooperatively provide optimal coverage in initial access, the MADRL algorithm should be developed to select the optimal SSB codebook for each cell, considering the UEs’ distribution and the capabilities of neighboring cells. One potential solution is to integrate the Graph Neural Network (GNN), which models heterogeneous NTNs as nodes with various features and helps to deal with large-scale and high-dimensional NTN scenarios, thereby facilitating the MADRL model training. Apart from that, the UEs associated with heterogeneous NTNs have distinct transmission power capabilities, which limit their connection to upper-tier NTNs such as satellite. Therefore, a lightweight model should be deployed at the UE to make the target NTN platform association decision based on local characteristics, including transmission power and battery.
  • Pilot Resource Allocation: In heterogeneous NTNs, complicated and overlapping coverage is a key cause of pilot interference. To mitigate this, the UE-level coordinated pilot resource allocation problem should be solved by the MADRL, where multiple cells learn to cooperatively minimize the channel correlation of UEs sharing the same pilot resource. However, UE-level allocation may increase the observation and action space significantly in MADRL, especially for higher-tier NTNs. Therefore, the hierarchical MADRL framework can be utilized to divide the coverage into several low-correlated smaller areas first, and then focus on the pilot resource allocation within each smaller area. To reduce the communication overhead during centralized training, channel charting can be a potential solution by extracting the low-dimensional features from the statistical channels [13].
  • Multi-hop IAB: Although the training of the routing optimization algorithm is delay-tolerant, the user data transmission may have stringent latency requirements. Therefore, multi-hop transmission of duplicate user data leads to increased transmission latency and channel capacity demands. To solve the problem, the hierarchical goal-oriented semantic communication can be integrated with the MADRL to reduce transmission overhead by prioritizing task-relevant information in each hop [14]. Additionally, the graph encoders, such as GNN, can also be incorporated into each agent’s observation, enhancing the adaptability of MADRL policies to topological changes.
  • Multi-timescale Control: In the multi-timescale trajectory optimization problem, the NTNs in different tiers experience different timescales of observation, action, and reward in MADRL due to varying altitude, propagation or processing delays, and control periods. Hence, a potential solution is to decouple the formulated hierarchical MADRL problem based on the timescales, and then optimize the subproblems iteratively [15]. Moreover, it is noted that goal-oriented semantic communication can also be utilized to quantify the importance of control signal transmission, which is closely coupled with the performance of the access links.
@@ IV-B Delay-sensitive Distributed Coordination
  • UL/DL Power Allocation: In this case, each NTN is required to deploy two heterogeneous agents to optimize the UL and DL power allocation, respectively. Specifically, the first agent detects the occurrence of CLI or RI based on the received signal and increases the uplink transmission power correspondingly. Then, the second agent optimizes the downlink transmission power based on the positions of neighboring cells to mitigate causing CLI or RI. It is noted that two agents of a single NTN work sequentially in the TDD frame structure, and same type of agents in different heterogeneous NTNs may work asynchronously due to dynamic TDD.
@@ V Case Study
In this section, to validate the effectiveness of the MADRL-empowered distributed coordination, we formulate a joint user scheduling and trajectory control optimization problem to maximize the average downlink throughput in a heterogeneous IAB-assisted NTN emergency communication scenario. Specifically, one T-UAV (IAB-Donor) and four U-UAVs (IAB-Node) are deployed at altitudes of 200 m and 100m, respectively. They serve the mobile ground users (G-UEs) through downlink transmission with high heterogeneity in terms of carrier frequency, bandwidth, and transmit power. During each time slot of 30 ms, we assume the incoming packets for each G-UE follow the Poisson process with a parameter of 4, and the latency for dropping packets is 10 time slots.
(a) Overall Average downlink throughput.
(b) Average downlink throughput of each UAV.
Figure 3: Average downlink throughput performance of the proposed TTS-MADDPG algorithm
To solve this problem, we propose a two-timescale MADDPG (TTS-MADDPG) algorithm under the CTDE framework. Specifically, this algorithm involves two groups of agents for two tasks: the first group (all UAVs) is responsible for the short-timescale user scheduling task at each time slot, while the second group (all U-UAVs) handles the long-timescale trajectory control task at every 5 time slots due to UAV’s mobility constraints. During the offline centralized training, global state information is utilized by the central critic to assist learning, while each agent’s actor is trained using only local observations.
The average downlink throughput results of our TTS-MADDPG algorithm during both training and testing phases are illustrated in Fig. 3, where we also include two benchmark schemes for the comparison: a round-robin (RR)-based scheduling method and a MADDPG-based scheduling approach. We conduct the simulation across 1000 episodes, with each episode consisting of 200 time slots, and the learning rates of actor and critic set to $10^{-4}$ and $10^{-3}$. In Fig. 3(a), for the converged performance, the TTS-MADDPG algorithm yields approximately 150% and 19% throughput gains over the RR-based and the MADDPG-based scheduling benchmarks, respectively, converging at around 175 Mbps. During testing, the TTS-MADDPG algorithm maintains stable performance that closely matches its converged performance during training, demonstrating strong generalization capability. In Fig. 3(b), with the TTS-MADDPG algorithm, T-UAV’s throughput declines slightly to 27.6 Mbps, as it learns to ensure sufficient scheduling on U-UAVs to serve the edge users and maximize the overall throughput. Moreover, each U-UAV achieves a noticeable throughput improvement to about 36 Mbps because of the optimal scheduling and real-time trajectory control.
The above results demonstrate that the proposed MADRL-based framework enables effective distributed coordination among heterogeneous UAV-based NTN platforms. While the case study focuses on the UAV-based scenario, it also exhibits rich heterogeneity in terms of physical characteristics, communication patterns, and the operation timescales, which acts as a solid foundation for extension to broader NTN architectures involving HAPs and satellites and for tackling other complicated challenges. Moreover, in the extended NTN scenario involving a vast number of agents and multiple timescales, it is challenging to achieve sufficient coordination, further leading to slow and unstable convergence in the MADRL training. To address these issues, except for the coordinated solutions mentioned above, the MADRL can be further integrated with the grouped training and decentralized execution framework or the graph encoders to reduce computational load and enhance the convergence.
@@ VI Conclusion
In this paper, we first summarize the unique characteristics of different NTN platforms and analyze their impact on the implementation of NTNs. We then focus on investigating the communication challenges for heterogeneous NTNs in individual layers, along with their potential coordinated solutions. Based on the emerging distributed algorithm, we further investigate the distributed MADRL algorithms tailored for each potential coordinated solution in heterogeneous NTNs, and analyze the CTDE deployment options according to their latency requirement. Importantly, we utilize the TTS-MADDPG algorithm to solve the joint trajectory design and user scheduling optimization problem in a heterogeneous NTN composed of both T-UAV and U-UAV, which validates the effectiveness of the distributed coordination algorithm among heterogeneous NTN platforms.
Jikang Deng is currently a Ph.D. student at King Abdullah University of Science and Technology (KAUST), Jeddah, Saudi Arabia. His research interests include wireless communication, non-terrestrial networks, and AI for communication.
Hui Zhou is currently an Assistant Professor at Coventry University. His research interests include non-terrestrial networks, integrated sensing and communication, unlicensed spectrum, machine learning, and semantic communication.
Mohamed-Slim Alouini is currently the Al-Khawarizmi Distinguished Professor of Electrical and Computer Engineering at King Abdullah University of Science and Technology (KAUST) and the holder of the UNESCO Chair on Education to Connect the Unconnected. He is a Fellow of the IEEE and OPTICA, and his research interests include the modeling, design, and performance analysis of wireless communication systems.